Find the following probabilities if the random variable X is normally distributed with


mean 15 and standard deviation 3.


(i) P(X<10)


(ii) P(10<X<25 )



Given that Z is a standard normal random variable, find z for each situation,


(i) The area to the left of Z is 0.2115


(ii) The area to the right of Z is 0.5000

1
Expert's answer
2022-04-28T10:13:42-0400

1.

(i)


"P(X<10)=P(Z<\dfrac{10-15}{3})\approx0.047790"

(ii)


"P(10<X<25)=P(Z<\dfrac{25-15}{3})"

"-P(Z\le\dfrac{10-15}{3})\approx0.9995709-0.0477904"

"\approx0.951781"

2.

(i)


"P(Z<z)=0.2115"


"z=-0.801227"

(ii)


"P(Z>z)=0.5000"


"z=0"